Homework: Section 4.5 Homework Save HW Score: 50%, 8 of 16 pts Question Help * Score: 0 of 1 pt 9 of 16 (8 complete 4.5.23 The table below displays results from exporiments with polygraph instruments Find the positive prediclive value for the test That is, Iind the probabilly that the subject lied, given that the test yields a positive result Did the Subject Actually Lie? No (Did Not Lie) Yes (Lied Positive test results Negative test results 12 30 45 The probability is Round to three decimal places as needed) ot Enter your answer in the answer box and then click Check Answer Clear All e (MAT 2018 P All parts showing
Answer #1

The positive predictive value (PPV) is defined as

{\text{PPV}}={\frac {\text{number of true positives}}{{\text{number of true positives}}+{\text{number of false positives}}}}={\frac {\text{number of true positives}}{\text{number of positive results}}}
Here number of true positive = 45

number of false positive = 12

Therefore number of positive calls = 45+12 = 57

So, PPV = \frac{45}{57} = 78.947%